OpenStudy (anonymous):

(2y^2 z^-3)/(5y^0z^5) simplify the expression

OpenStudy (lgbasallote):

\[\huge \frac{2y^2 z^{-3}}{5y^0 z^5} \implies \frac 25 \times \frac{y^2}{y^0} \times \frac{z^{-3}}{z^5}\] does that help?